Find the value of p and the value of q
x^-6x+15 = (x-p)^+q

Answer by stanbon(75887)

Find the value of p and the value of q
x^2-6x+15 = (x-p)^2+q
---------------
x^2 - 6x + 15 = x^2 - 2px + p^2 + q
-----------------------------------------
Equating the corresponding coefficients you get:
-2p = -6
p^2+q = 15
---------------
solve the 1st for "p":
p = 3
----
Substitute into the 2nd equation and solve for "q":
9 + q = 15
q = 6
==================
Cheers,
Stan H.